# Spatio-Temporal Context Prompting for Zero-Shot Action Detection

### v2

@@ 1 Introduction
  • We propose a complete benchmark on J-HMDB, UCF101-24 and AVA datasets to evaluate performance on Zero-Shot Spatio-Temporal Action Detection. The experiments demonstrate the strong generalization capabilities of our method, and show the ability to individually detect unseen actions within the same video.
@@ 2 Related Work
Spatio-Temporal Action Detection. Typical action detection methods mostly use the two-stage pipeline, which means first localizing people in a video, and then performing action classification based on the features of these people. Most of these methods utilize additional person detectors like Faster R-CNN [21] to generate actor bounding boxes, which are used to perform RoIAlign [8] on the video features generated by the 3D CNN backbone to obtain the person features. While [5] directly utilizes naive actor features to classify actions, [4, 24, 19] further exploit relation modeling to combine more information about human and environmental interactions. Besides, some methods [23, 31, 6, 1] optimize the two-stage networks by a joint loss in an end-to-end framework. Recently, there are query-based detectors that simultaneously predict the actor positions and their corresponding actions in a single-stage manner. [35] utilizes a set of tubelet-queries to sample features from video representation, while [29] samples discriminative features based on the multi-scale feature map. Although the above works have made significant progress in traditional fully-supervised scenarios, our goal is to better recognize unseen actions through zero-shot learning.
@@ 3.2 Overall Architecture
Subsequently, to fully leverage CLIP’s visual knowledge, we jointly input person and context tokens into the image encoder. The Multi-Head Self Attention (MHSA) in each encoder layer is employed to guide us in achieving the following three objectives: (i) performing further spatial modeling on the input context tokens to obtain spatial-temporal tokens. (ii) modeling person-person and person-context interaction through the mutual influence between tokens. (iii) identifying the interest tokens most relevant to each person’s actions through attention weight. On the textual side, we initially utilize the CLIP text encoder to generate the original text features for class names. Then, followed by each image encoder layer, the Context Prompting layer will use context tokens to prompt each label. In this process, considering that the videos in J-HMDB and UCF101-24 only contain single action, we can treat all context tokens as relevant to this action. Hence, we use all context tokens for prompting, resulting in everyone in the same frame sharing the same text features. However, in AVA, to discern different actions by multiple individuals, we utilize context tokens that each person deems important (i.e., interest tokens) to prompt their respective text features. Finally, we use the person tokens and label features output by the last image encoder layer and Context Prompting layer to calculate the cosine similarities, which are used as the action classification scores.
In the training process, in order to retain the generalization capability of CLIP for zero-shot tasks, we freeze the pretrained weights in the image and text encoders, and only train our additional learnable modules. Besides, we insert the LoRA trainable matrices into the Feed-Forward Network (FFN) of each image encoder layer, which can further adapt CLIP model to detect actions without affecting its well-aligned visual-language features.
@@ 3.3 Person-Context Interaction
where $e^{temp}$ is the temporal encoding and $LN$ stands for layer normalization. After temporal modeling, we can obtain context tokens $\tilde{Z_{i}}\in\mathbb{R}^{D}$ at each position $i\in\{1,\dots,N\}$ that have aggregated temporal information, which will be used to model the interaction with person.
@@ 3.4 Interest Token Spotting
In a multi-action dataset like AVA, the same timestamp may encompass various actions performed by multiple individuals, and all the context tokens will contain information about different actions. Therefore, we introduce Interest Token Spotting, a mechanism that employs a personal perspective to extract context tokens most relevant to each individual’s actions. Subsequently, we utilize these tokens for prompting to generate personalized text features. In this process, we also leverage CLIP’s pretrained visual knowledge to find each person’s interest tokens. To be more specific, we exploit the attention weight calculated by the MHSA in each image encoder layer as an indicator of the token importance. The Multi-Head Self Attention will first calculate an attention map $M_{i}\in\mathbb{R}^{C\times{C}}$ based on the query and key of each head, where $i\in\{1,\dots,num\_heads\}$, and $C$ is the number of input tokens. Then, we average the attention maps of each head to obtain an importance score matrix $M\in\mathbb{R}^{C\times{C}}$. In this matrix, each row represents the importance of every token to a certain token. For instance, $M(i,j)$ represents how important token $j$ is to token $i$. Based on this, we use a person’s token as the row index, and select the top $K$ highest among all $C$ importance scores. These selected $K$ indexes are the token positions that the person is interested in. After that, we pass each token through the MHSA and FFN of this encoder layer, and use the selected $K$ indexes to obtain the interest tokens.
@@ 3.5 Context Prompting
The Context Prompting layer primarily consists of Cross-Attention (CA), where text features serve as the query, and the context tokens act as key and value. This setup facilitates the gradual absorption of visual information into the text content. In AVA, we further narrow down the scope of the extracted information for each person to their personal interest tokens. Given an image with $B$ detected individuals, we need to classify their actions into $N_{L}$ possible labels. First, we assign the same set of original text features which are obtained by the CLIP text encoder to these $B$ people. Subsequently, we employ a linear projector to project each person’s interest tokens to the same dimension as the text features. Following this, both the interest tokens and the text features are sent to the Context Prompting layer for prompting. The following equations describe how it works:
@@ 4.3 ZS-AVA
AVA[7] is a large-scale action detection dataset which contains multiple actions within a single video. It consists of 235 videos for training and 64 videos for validation. Each video lasts for 15 minutes and is annotated by sampling one keyframe per second. For AVA, since the same video contains multiple actions, and some action categories exist in multiple videos, it becomes challenging to find a sufficient amount of training data and testing data with disjoint labels. Instead, we randomly select some training videos, ensuring that they all lack samples of the same classes. These missing classes are then treated as unseen classes for evaluation. During the evaluation phase, we test all classes in the validation videos, but the focus is solely on evaluating the performance of unseen classes. Under this setting, we propose three splits. When the amounts of training and testing data for these splits are nearly the same, we select different combinations of three action types — pose action, object interaction, and person interaction as unseen classes, allowing for a more comprehensive evaluation.
@@ 5.1 Experimental Setup
Person Detector: In the following experiments, we employ Faster R-CNN [13] with a ResNet-50-FPN [14] backbone pretrained on MSCOCO [15] for person detection. For J-HMDB and AVA, we directly inference on the test data with pretrained person detector. While for UCF101-24, since the images have lower resolution, we further use the ground-truth of training classes to finetune the person detector for 10 epochs in each label split. Besides, to remove the false positives, we keep the detected box with the highest confident score $S$ in each frame, then we select the boxes with scores higher than $S-T$ from the rest, where $T$ is 0.001 for J-HMDB and 0.7 for both UCF101-24 and AVA.
Hyperparameters: For all the experiments on J-HMDB and UCF101-24, we employ ViT-B/16 as our CLIP backbone and use the same hyper-parameters as follows: Training for 3K iterations with batch size of 8. We use SGD as our optimizer and the base learning rate is set to 2.5e-4. As for AVA, we use ViT-L/14 backbone and train the model for 20K iterations with a base learning rate of 4e-4.
@@ 5.2 Zero-Shot Spatial-Temporal Action Detection
Frame mAP@0.5 Method split 1 split 2 split 3 avg avg* baseline 8.67 8.57 2.77 6.67 10.09 baseline (person crop) 8.22 5.05 3.04 5.44 7.55 iCLIP [9] 4.04 9.08 1.91 5.01 6.91 Vita-CLIP [28] 4.25 4.29 0.64 3.06 10.45 ST-CLIP (Ours) 12.85 10.17 4.01 9.01 11.76
@@
Table 1: Evaluation on ZS-AVA. * denotes using the ground-truth bounding boxes of the test data. All methods employ ViT-L/14 backbone. In the inference stage of Vita-CLIP [28], we use two different tracklets: (1) tracklets obtained by associating detected boxes with ByteTrack [34], and (2) tracklets with ground-truth boxes provided by AVA official.
@@
In addition to using detected person boxes to measure the actual results on zero-shot spatial-temporal action detection, we also provide results using ground-truth bounding boxes of the test data. This allows us to analyze the performance of all methods without the influence of localization errors. The results of each label split using ground-truth bounding boxes are provided in the supplementary.
Table 1 shows our results on detecting unseen actions in AVA. The main distinction between our method and the video classification approaches is our capability to detect different unseen actions within the same video, which are difficult to achieve. Firstly, in the training process of those methods, providing a fixed video label is challenging due to the presence of numerous different actions. Additionally, during the inference stage, these methods tend to detect that everyone in the video has the same action. To study the feasibility of the video classification method, we further narrowed the scope of classification from the entire video to tracklets to avoid misclassifying different actions into the same category. Besides, considering the limited prior research in this area, we follow [9] to implement a naive baseline for comparison with our method. For a frame with detected individuals, the baseline utilizes the pretrained image encoder of CLIP to extract the image feature of this frame. Subsequently, it calculates the cosine similarities with the text features of each class name, which are then considered as the action classification scores for these individuals. Since the baseline regards people in the same frame as having the same actions, we also implement the person crop method. This involves cropping out parts of each person to obtain their respective image features for classification. Nevertheless, this cropping method may result in reduced performance for the baseline in split 1 and split 2, as it captures insufficient information. The results show that our ST-CLIP achieves the best average performance, whether using detected boxes or ground-truth boxes, demonstrating its effectiveness in detecting individual unseen actions.
@@
We present the experimental results on ZS-JHMDB and ZS-UCF in Tabs. 2 and 3. Firstly, since each video in both datasets contains only a single action, a straightforward approach is to conduct action detection through zero-shot video classification. These methods can initially classify the entire video into an action class and then consider all detected individuals in the video as performing this action. However, since our method concentrates on detecting different actions for each person, individuals in the same video may be classified into different actions. This general setting can result in other video classification methods having an advantage over us in these two datasets. For a fair comparison with the other methods, we further adopt the assumption that a video contains only one action. In this context, we perform soft voting on each person’s classification score, extending our method to suit this scenario.
We first present the results on ZS-JHMDB in Table 2. Regardless of whether the assumption is applied, our ST-CLIP outperforms others in most label splits, and thus achieves the best average performance, along with 3.82 mAP higher than the state-of-the-art method [28]. Additionally, by using ground-truth bounding boxes to eliminate localization errors, our average performance improves to 90.12 mAP, which is 4.1 mAP higher than [28].
Table 3 presents the results on ZS-UCF. Firstly, without the assumption, our method exhibits a 2.31 mAP improvement over [9], which also focuses on detecting individual actions. It is worth mentioning that the localization errors have a noticeable impact on our performance in this dataset. Since there are instances where irrelevant people who are not performing actions are detected, these individuals should be considered as part of the background. However, these false positive cases will also contribute to the soft voting process, leading to our classification results being slightly inferior to other methods that solely rely on sampled frames to determine video labels. In this case, our method still outperforms [27, 12], and exhibits similar performance to [18]. With ground-truth bounding boxes, our method demonstrates the second-best average performance, surpassing [18] and achieving results comparable to [28]. The results on ZS-JHMDB and ZS-UCF show that our ST-CLIP can perform competitively with other video classification methods when processing single-action videos.
@@
Frame mAP@0.5 Method split 1 split 2 split 3 split 4 avg avg* Without the assumption of single-action video baseline 64.63 71.04 82.13 75.88 73.42 81.21 iCLIP [9] 66.53 69.99 82.88 71.84 72.81 79.01 ST-CLIP (Ours) 74.55 74.97 83.59 83.22 79.08 85.53 With the assumption of single-action video ActionCLIP [27] 69.18 75.28 77.11 76.55 74.53 82.07 A5 [12] 50.92 66.06 69.07 60.21 61.57 67.39 X-CLIP [18] 72.91 72.62 80.02 77.78 75.83 83.11 Vita-CLIP [28] 68.60 82.35 84.95 80.19 79.02 86.02 ST-CLIP (Ours) 79.62 78.70 85.84 87.19 82.84 90.12
@@
Table 2: Evaluation on ZS-JHMDB. * denotes using the ground-truth boxes of the test data. All methods use ViT-B/16 backbone.
@@
Frame mAP@0.5 Method split 1 split 2 split 3 split 4 avg avg* Without the assumption of single-action video baseline 48.37 52.84 39.76 51.92 48.22 90.49 iCLIP [9] 50.34 52.75 39.73 47.95 47.69 85.47 ST-CLIP (Ours) 49.09 54.95 42.05 53.92 50.00 91.80 With the assumption of single-action video ActionCLIP [27] 52.64 55.01 38.04 51.84 49.38 89.72 A5 [12] 46.78 47.36 41.23 53.18 47.14 85.60 X-CLIP [18] 50.10 56.94 44.52 55.35 51.73 94.12 Vita-CLIP [28] 52.52 57.22 45.12 57.32 53.05 96.57 ST-CLIP (Ours) 51.36 56.30 43.12 54.52 51.33 95.02
@@
Table 3: Evaluation on ZS-UCF. * denotes using the ground-truth boxes of the test data. All methods use ViT-B/16 backbone.

### v3

@@ 1 Introduction
  • We propose a complete benchmark on J-HMDB, UCF101-24, and AVA datasets to evaluate performance on Zero-Shot Spatio-Temporal Action Detection. The experiments demonstrate the strong generalization capabilities of our method, and show the ability to individually detect unseen actions within the same video.
@@ 2 Related Work
Spatio-Temporal Action Detection. Typical action detection methods mostly use the two-stage pipeline, which means first localizing people in a video, and then performing action classification based on the features of these people. Most of these methods utilize additional person detectors like Faster R-CNN [22] to generate actor bounding boxes, which are used to perform RoIAlign [9] on the video features generated by the 3D CNN backbone to obtain the person features. While [6] directly utilizes naive actor features to classify actions, [5, 26, 20] further exploit relation modeling to combine more information about human and environmental interactions. Besides, some methods [25, 33, 7, 2] optimize the two-stage networks by a joint loss in an end-to-end framework. There are also several methods designed to handle similar tasks in unsupervised scenarios. The goal of [23] is to localize the time and space of actions in the video without using bounding box annotations during training. [1] propose a domain adaptation framework for action detection; however, their approach is limited to detecting fixed action classes, whereas our goal is to operate in an open-vocabulary setting.
@@ 3.2 Overall Architecture
Subsequently, to fully leverage CLIP’s visual knowledge, we jointly input person and context tokens into the image encoder. The Multi-Head Self Attention (MHSA) in each encoder layer is employed to guide us in achieving the following three objectives: (i) performing further spatial modeling on the input context tokens to obtain spatial-temporal tokens. (ii) modeling person-person and person-context interaction through the mutual influence between tokens. (iii) identifying the interest tokens most relevant to each person’s actions through attention weight. On the textual side, we initially utilize the CLIP text encoder to generate the original text features for class names. Then, followed by each image encoder layer, the Context Prompting layer will use context tokens to prompt each label. In this process, considering that the videos in J-HMDB and UCF101-24 only contain a single action, we can treat all context tokens as relevant to this action. Hence, we use all context tokens for prompting, resulting in everyone in the same frame sharing the same text features. However, in AVA, to discern different actions by multiple individuals, we utilize context tokens that each person deems important (i.e., interest tokens) to prompt their respective text features. Finally, we use the person tokens and label features output by the last image encoder layer and Context Prompting layer to calculate the cosine similarities, which are used as the action classification scores.
In the training process, in order to retain the generalization capability of CLIP for zero-shot tasks, we freeze the pretrained weights in the image and text encoders, and only train our additional learnable modules. Besides, we insert the LoRA trainable matrices into the Feed-Forward Network (FFN) of each image encoder layer, which can further adapt the CLIP model to detect actions without affecting its well-aligned visual-language features.
@@ 3.3 Person-Context Interaction
where $e^{temp}$ is the temporal encoding and $LN$ stands for layer normalization. After temporal modeling, we can obtain context tokens $\tilde{Z_{i}}\in\mathbb{R}^{D}$ at each position $i\in\{1,\dots,N\}$ that have aggregated temporal information, which will be used to model the interaction with a person.
@@ 3.4 Interest Token Spotting
In a multi-action dataset like AVA, the same timestamp may encompass various actions performed by multiple individuals, and all the context tokens will contain information about different actions. Therefore, we introduce Interest Token Spotting, a mechanism that employs a personal perspective to extract context tokens most relevant to each individual’s actions. Subsequently, we utilize these tokens for prompting to generate personalized text features. In this process, we also leverage CLIP’s pretrained visual knowledge to find each person’s interest tokens. To be more specific, we exploit the attention weight calculated by the MHSA in each image encoder layer as an indicator of the token importance. The Multi-Head Self Attention will first calculate an attention map $M_{i}\in\mathbb{R}^{C\times{C}}$ based on the query and key of each head, where $i\in\{1,\dots,num\_heads\}$, and $C$ is the total number of person tokens and context tokens. Then, we average the attention maps of each head to obtain an importance score matrix $M\in\mathbb{R}^{C\times{C}}$. In this matrix, each row represents the importance of every token to a certain token. For instance, $M(i,j)$ represents how important token $j$ is to token $i$. Based on this, we use a person’s token as the row index, and select the top $K$ highest among all $C$ importance scores. These selected $K$ indexes are the token positions that the person is interested in. After that, we pass each token through the MHSA and FFN of this encoder layer, and use the selected $K$ indexes to obtain the interest tokens.
@@ 3.5 Context Prompting
The Context Prompting layer primarily consists of Cross-Attention (CA), where text features serve as the query, and the context tokens act as key and value. This setup facilitates the gradual absorption of visual information into the text content. In AVA, we further narrow down the scope of the extracted information for each person to their personal interest tokens. Given an image with $B$ detected individuals, we need to classify their actions into $N_{L}$ possible labels. First, we assign the same set of original text features that are obtained by the CLIP text encoder to these $B$ people. Subsequently, we employ a linear projector to project each person’s interest tokens to the same dimension as the text features. Following this, both the interest tokens and the text features are sent to the Context Prompting layer for prompting. The following equations describe how it works:
@@ 4.3 ZS-AVA
AVA[8] is a large-scale action detection dataset that contains multiple actions within a single video. It consists of 235 videos for training and 64 videos for validation. Each video lasts for 15 minutes and is annotated by sampling one keyframe per second. For AVA, since the same video contains multiple actions, and some action categories exist in multiple videos, it becomes challenging to find a sufficient amount of training data and testing data with disjoint labels. Instead, we randomly select some training videos, ensuring that they all lack samples of the same classes. These missing classes are then treated as unseen classes for evaluation. During the evaluation phase, we test all classes in the validation videos, but the focus is solely on evaluating the performance of unseen classes. Under this setting, we propose three splits. When the amounts of training and testing data for these splits are nearly the same, we select different combinations of three action types — pose action, object interaction, and person interaction as unseen classes, allowing for a more comprehensive evaluation.
@@ 5.1 Experimental Setup
Person Detector: In the following experiments, we employ Faster R-CNN [14] with a ResNet-50-FPN [15] backbone pretrained on MSCOCO [16] for person detection. For J-HMDB and AVA, we directly inference on the test data with pretrained person detector. For UCF101-24, since the images have lower resolution, we further use the ground truth of training classes to finetune the person detector for 10 epochs in each label split. Besides, to remove the false positives, we keep the detected box with the highest confident score $S$ in each frame, then we select the boxes with scores higher than $S-T$ from the rest, where $T$ is 0.001 for J-HMDB and 0.7 for both UCF101-24 and AVA.
Hyperparameters: For all the experiments on J-HMDB and UCF101-24, we employ ViT-B/16 as our CLIP backbone and use the same hyper-parameters as follows: Training for 3K iterations with a batch size of 8. We use SGD as our optimizer and the base learning rate is set to 2.5e-4. As for AVA, we use the ViT-L/14 backbone and train the model for 20K iterations with a base learning rate of 4e-4.
5.2 Compared Methods
In addition to comparing our approach with iCLIP [10], which addresses the same task as ours, we also evaluate it against the following methods.
Baseline: we follow [10] to implement a naive baseline for comparison. For a frame with detected individuals, the baseline utilizes the pretrained image encoder of CLIP to extract the image feature of this frame. Subsequently, it calculates the cosine similarities with the text features of each class name, which are then considered as the action classification scores for these individuals. Since the baseline regards people in the same frame as having the same actions, we also implement the person crop method for multi-action videos. This involves cropping out parts of each person to obtain their respective image features for classification.
ViCLIP [30]: We also experiment with a video-language model for individual action classification. First, we extract the video feature map using a video encoder and apply average pooling along the temporal dimension. Then, for each person, we use their bounding box to perform ROIAlign and max pooling on the feature map, obtaining a person-specific feature for classification.
Video classification methods [29, 13, 19, 31]: since each video in both ZS-JHMDB and ZS-UCF contains only a single action, a straightforward approach is to conduct action detection through video classification. These methods can initially classify the entire video into an action class and then consider all detected individuals in the video as performing this action. As for ZS-AVA, the main distinction between our method and these approaches is our capability to detect different actions within the same video, which is difficult to achieve. Firstly, in the training process of these methods, providing a fixed video label is challenging due to the presence of numerous different actions. Additionally, during the inference stage, these methods tend to detect that everyone in the video has the same action. To study the feasibility of these methods, we further narrow the scope of classification from the entire video to tracklets to avoid misclassifying different actions into the same category.
@@ 5.3 Zero-Shot Spatial-Temporal Action Detection
Frame mAP@0.5 Method split 1 split 2 split 3 avg H avg* H* baseline 8.67 8.57 2.77 6.67 7.38 10.09 10.58 baseline (person crop) 8.22 5.05 3.04 5.44 6.42 7.55 8.63 iCLIP [10] 4.04 9.08 1.91 5.01 7.59 6.91 10.37 Vita-CLIP [31] 4.25 4.29 0.64 3.06 4.49 10.45 13.94 ST-CLIP (Ours) 12.85 10.17 4.01 9.01 11.41 11.76 15.05
@@
Table 1: Evaluation on ZS-AVA. * denotes using the ground-truth bounding boxes of the test data. All methods employ the ViT-L/14 backbone. In the inference stage of Vita-CLIP [31], we use two different tracklets: (1) tracklets obtained by associating detected boxes with ByteTrack [36], and (2) tracklets with ground-truth boxes provided by AVA official.
@@
In this section, we present the experimental results for unseen classes in each label split. We also calculate the harmonic mean (H) of the average performance of both base and unseen classes, to provide a more comprehensive assessment. In addition to using detected person boxes to measure the actual results on zero-shot spatial-temporal action detection, we also provide results using ground-truth bounding boxes of the test data. This allows us to analyze the performance without the influence of localization errors.
Table 1 shows our results on detecting unseen actions in AVA. First of all, the baseline method aligns image features during both the pretraining phase of CLIP and the inference phase of action detection, leading to good performance on unseen classes. It is worth noting that although iCLIP [10] and the baseline method achieve similar harmonic mean, iCLIP’s performance on unseen classes is relatively poor. This suggests that their interaction modeling affects generalization ability. In contrast, our Person-Context Interaction approach leverages CLIP’s pretrained knowledge, resulting in superior performance on both base and unseen classes.
@@
We present the experimental results on ZS-JHMDB and ZS-UCF in Tabs. 2 and 3. Since our method focuses on detecting different actions for each person, individuals in the same video may be classified into different actions. This general setting can result in other video classification methods having an advantage over us in these two datasets. For a fair comparison with the other methods, we further adopt the assumption that a video contains only one action. In this context, we perform soft voting on each person’s classification score, extending our method to suit this scenario.
We first present the results on ZS-JHMDB in Table 2. Firstly, the video-language model [30] with ROI align has difficulty accurately detecting unseen actions. We speculate that this is primarily due to the alignment of global video features (CLS token) during the pretraining process of ViCLIP. As a result, the person-specific features extracted from the video feature map could not align well with the text features. Without the assumption, our ST-CLIP achieves the highest average performance on unseen classes, along with the best harmonic mean. With the assumption applied, our method still outperforms other video classification approaches, achieving an average performance 3.82 mAP higher than [31]. Furthermore, by using ground-truth bounding boxes to eliminate localization errors, our performance improves further to 90.12 mAP.
@@
Frame mAP@0.5 Method split 1 split 2 split 3 split 4 avg H avg* H* Without the assumption of single-action video baseline 64.63 71.04 82.13 75.88 73.42 65.45 81.21 71.82 ViCLIP [30] 52.29 72.29 53.59 70.58 62.19 68.03 68.15 74.05 iCLIP [10] 66.53 69.99 82.88 71.84 72.81 74.47 79.01 81.09 ST-CLIP (Ours) 74.55 74.97 83.59 83.22 79.08 77.65 85.53 84.32 With the assumption of single-action video ActionCLIP [29] 69.18 75.28 77.11 76.55 74.53 75.82 82.07 83.25 A5 [13] 50.92 66.06 69.07 60.21 61.57 68.94 67.39 75.34 X-CLIP [19] 72.91 72.62 80.02 77.78 75.83 77.13 83.11 84.34 Vita-CLIP [31] 68.60 82.35 84.95 80.19 79.02 80.46 86.02 87.49 ST-CLIP (Ours) 79.62 78.70 85.84 87.19 82.84 80.96 90.12 88.48
@@
Table 2: Evaluation on ZS-JHMDB. * denotes using the ground-truth boxes of the test data. All methods use the ViT-B/16 backbone.
@@
Frame mAP@0.5 Method split 1 split 2 split 3 split 4 avg H avg* H* Without the assumption of single-action video baseline 48.37 52.84 39.76 51.92 48.22 49.00 90.49 82.45 ViCLIP [30] 41.37 43.93 24.14 34.04 35.87 43.72 63.07 72.50 iCLIP [10] 50.34 52.75 39.73 47.95 47.69 54.46 85.47 90.35 ST-CLIP (Ours) 49.09 54.95 42.05 53.92 50.00 54.41 91.80 91.34 With the assumption of single-action video ActionCLIP [29] 52.64 55.01 38.04 51.84 49.38 55.36 89.72 92.91 A5 [13] 46.78 47.36 41.23 53.18 47.14 54.24 85.60 90.86 X-CLIP [19] 50.10 56.94 44.52 55.35 51.73 56.96 94.12 95.30 Vita-CLIP [31] 52.52 57.22 45.12 57.32 53.05 57.65 96.57 95.94 ST-CLIP (Ours) 51.36 56.30 43.12 54.52 51.33 55.94 95.02 94.75
@@
Table 3: Evaluation on ZS-UCF. * denotes using the ground-truth boxes of the test data. All methods use the ViT-B/16 backbone.
Table 3 presents the results on ZS-UCF. It is worth mentioning that the localization errors have a noticeable impact on our performance in this dataset. Since there are instances where irrelevant people who are not performing actions are detected, these individuals should be considered as part of the background. However, these false positive cases will also contribute to the soft voting process, leading to our classification results being slightly inferior to other methods that solely rely on sampled frames to determine video labels. In this case, our method still outperforms [29, 13], and exhibits similar performance to [19]. With ground-truth bounding boxes, our method achieves the second-best average performance. However, it is important to note that the UCF101-24 dataset has relatively low image quality, with most characters occupying only a small portion of the frame. This impacts the accuracy of our individual classification, putting us at a slight disadvantage compared to [31], which relies on the entire sampled frame.
5.4 Ablation Study
We present an ablation study in Tabs. 4, 6, 6, 8 and 8 to investigate different design choices in our method. The experiments are performed on the label split 1 of ZS-JHMDB and ZS-AVA.
Proposed components: We first investigate the importance of each component in Table 4. On J-HMDB, our proposed Person-Context Interaction effectively models the relationship between individuals and their surroundings, resulting in a 2.35 mAP improvement compared to the baseline. Furthermore, the Context Prompting module leverages context information to enhance text content, leading to an additional improvement of 7.57 mAP. On AVA, the above two components also demonstrate their effectiveness. Additionally, in multi-action videos, our Interest Token Spotting can identify context tokens most relevant to individual actions for prompting, further enhancing performance.